# AlpacaFarm: A Simulation Framework for Methods that Learn from Human Feedback

### 06 Dec 2023 v3

@@ Abstract
Large language models (LLMs) such as ChatGPT have seen widespread adoption due to their ability to follow user instructions well. Developing these LLMs involves a complex yet poorly understood workflow requiring training with human feedback. Replicating and understanding this instruction-following process faces three major challenges: the high cost of data collection, the lack of trustworthy evaluation, and the absence of reference method implementations. We address these challenges with AlpacaFarm, a simulator that enables research and development for learning from feedback at a low cost. First, we design LLM prompts to simulate human feedback that are 45x cheaper than crowdworkers and display high agreement with humans. Second, we propose an automatic evaluation and validate it against human instructions obtained on real-world interactions. Third, we contribute reference implementations for several methods (PPO, DPO, best-of-$n$, expert iteration, and more) that learn from pairwise feedback. Finally, as an end-to-end validation of AlpacaFarm, we train and evaluate eleven models on 10k pairs of real human feedback and show that the rankings of models trained in AlpacaFarm match the rankings of models trained on human data. As a demonstration of the research possible in AlpacaFarm, we find that methods that use a reward model can substantially improve over supervised fine-tuning and that our reference PPO implementation leads to a +10% improvement in win-rate against Davinci003. We release all components of AlpacaFarm at https://github.com/tatsu-lab/alpaca_farm.
Figure 1: AlpacaFarm is a simulation sandbox that enables fast and cheap experimentation on methods that learn from human feedback. It simulates human feedback with API LLMs, provides a validated evaluation protocol, and offers a suite of reference method implementations. Researchers can rapidly iterate on model development and transfer their methods to training on human data to maximize performance.
@@ 1 Introduction
Large language models (LLMs) (Bommasani21, ; Brown20, ; OpenAI23, ) have demonstrated unprecedented capabilities in following diverse and open-ended instructions (Ouyang22, ; Askell21, ; Longpre23, ). These achievements have often been attributed to the fine-tuning of pretrained LLMs using human feedback, but this process remains poorly understood due to the lack of published information on the training methods from LLM vendors. For example, it was recently revealed that only the Davinci003 model in the instruct series of OpenAI models used reinforcement learning (RL) with the PPO algorithm OpenAI*a , leading some to question the importance of RL in the training process. Understanding and improving these methods requires open and transparent replications of the training process, but this remains challenging due to the cost and complexity associated with methods that learn from human feedback.
@@
Our goal is to facilitate research and development on instruction following models and methods that learn from human feedback. We identify three main challenges: the high cost of data annotation, the lack of automated evaluation for model development, and the absence of validated implementations of existing methods. To address these three challenges, we introduce AlpacaFarm (Figure 1), a simulation sandbox that enables experimentation at a low cost. Using AlpacaFarm, researchers can rapidly iterate on method development in simulation and transfer these insights to build high-performance systems with actual human feedback.
@@
For the first challenge of data annotation costs, AlpacaFarm simulates human annotators with API LLMs that are faster and lower cost. To collect simulated feedback data, we design prompts for API LLMs (e.g. GPT-4) that enable us to simulate human pairwise comparisons at a cost that is 45x cheaper than crowdworkers, and tune these prompts to faithfully capture many aspects of human annotators, such as their quality judgments, inter-annotator variability, and stylistic preferences.
@@
For the second challenge of automated evaluation, we design an automatic evaluation protocol that attempts to quantify system performance on simple but realistic real-world human instructions. Improving evaluations of open-ended instruction following has been challenging due to the cost and non-replicability of human evaluation, the lack of real human interaction data, and the diversity of natural human instructions. To address this, we use instructions from user interactions (Alpaca Demo (Taori23, )) as a reference for simple but real human interactions and show that we can combine existing public evaluation datasets to mimic this evaluation. Quantitative evaluations of system rankings on our evaluation data show a high correlation with system rankings on the Alpaca Demo instructions.
@@
For the third challenge of missing reference implementations, we implement and test several popular learning algorithms including PPO (Schulman17, ), expert iteration (Anthony17, ), and Quark Lu22*b , and release reference implementations. We show that, among the methods we studied, PPO with a surrogate reward model is the most effective training-time method in our leaderboard, improving the win-rate against Davinci003 of an instruction fine-tuned LLaMA 7B model from 44% to 55%. Other baselines that have been validated on simpler tasks fall short in comparison, highlighting the importance of testing these algorithms in a real instruction-following setting.
@@
As an end-to-end evaluation of the AlpacaFarm, we compare eleven methods trained and evaluated in AlpacaFarm with the same methods trained and evaluated on actual human feedback. We show that the method rankings obtained from developing on AlpacaFarm closely agree with the method rankings obtained from training on actual human data (Spearman correlation of 0.98) and that the best method in AlpacaFarm leads to substantial gains with human feedback. Finally, we find that AlpacaFarm can replicate qualitative behaviors of human feedback such as over-optimization of the reward model, suggesting that AlpacaFarm serves as an effective way for researchers to rapidly study and develop methods that learn from human feedback.
@@ 2 Background & problem statement
To begin, we introduce the instruction following task and the pairwise-comparison-based human feedback setting that we study. With this background, we formally define the goals of developing a low-cost simulator for studying instruction following and learning from human feedback.
@@ 2.1 Learning to follow instructions
While there are a rich set of methods that learn from human feedback to directly optimize $R$ (see Section 6), in this work we focus on the setting of learning from pairwise feedback (LPF) due to its central role in recent instruction-following LLMs Ouyang22 . The starting point of this process is a model that is fine-tuned on instruction-following demonstrations $(x,y)$, which we denote as $p^{\text{SFT}}_{\theta}(y\mid x)$. The LPF process then involves taking pairs of samples from $p^{\text{SFT}}_{\theta}$, querying humans for which sample within each pair is better, and learning from this pairwise feedback. Since all methods start from the SFT base, we use $p_{\theta}$ for notational simplicity.
@@
Once instruction-following models are trained, researchers need to evaluate these models. One common approach for evaluating models is pairwise model evaluation (Chiang23, ; Geng23, ; Bai22, ), which performs pairwise comparisons on outputs generated by the model $p_{\theta}$ and a reference model $p_{\text{ref}}$. Concretely, we collect pairwise preference for two models ($\{(x^{(j)},y_{\theta}^{(j)},y_{\text{ref}}^{(j)},z^{(j)})\}_{j}$), which is aggregated by computing the average win-rate – the percentage of times $p_{\theta}$ is preferred to $p_{\text{ref}}$. Researchers can then compare LPF models by their win-rates against the same reference model.
@@ 2.2 Problem statement
For pairwise feedback, we substitute human preference judgements $z_{\text{human}}\sim p_{\text{human}}(z\mid x,y_{0},y_{1})$ with a simulated preference $z_{\text{sim}}\sim p_{\text{sim}}(z\mid x,y_{0},y_{1})$ using API LLMs. Our goal is to construct a $p_{\text{sim}}$ that is both low-cost and faithfully captures different aspects of human preference feedback, such as quality judgments, inter-annotator agreement rates, and stylistic preferences.
For evaluations, we evaluate system outputs using the pairwise preference simulator and identify evaluation datasets that reflect natural human-LLM interactions. The goal for our evaluations is to ensure that system rankings on the new evaluation dataset closely match both human rankings and rankings on instructions from real usage of the Alpaca Demo.
@@
AlpacaFarm combines these three components into a simulation framework for learning from pairwise feedback. We evaluate the complete system by an end-to-end workflow of developing methods in simulation and transferring the insights to the real world.
Concretely, we will run each method $M$ on the simulated preferences (called $M_{\text{sim}}$) and evaluate with simulated rankings $p_{\text{sim}}$. In parallel, we will run $M$ on human preferences (called $M_{\text{human}}$) and evaluate with human rankings $p_{\text{human}}$. We consider AlpacaFarm to be successful if the simulated method rankings correlate well with the human method rankings. The rest of this work will present the details of the pairwise feedback and evaluation design (Section 3), evaluate these designs (Section 4), and analyze the different reference methods we implemented in the AlpacaFarm (Section 5).
@@ 3.2 Designing simulated pairwise preference $p_{\text{sim}}$
Equipped with the Alpaca instruction data, we now describe the design of our simulated annotator for pairwise preferences. Our core proposal is to design the simulator $p_{\text{sim}}(z\mid x,y_{0},y_{1})$ by prompting OpenAI API LLMs. While using LLMs as a proxy for annotators has become increasingly popular (Chiang23, ; Liu23*a, ), using LLMs as part of a simulation environment poses major additional challenges. Our simulated preferences $p_{\text{sim}}$ must not only have a high agreement with human preferences $p_{\text{human}}$, it must also capture other qualitative aspects of human feedback such as inter- and intra-annotator inconsistencies. Intuitively, the noise and variability in pairwise feedback are key parts of the challenge in the LPF problem and we find that ignoring these factors leads to a simulator that diverges substantially from real-world behavior (Section 4.3).
@@
To start with, we design prompts by providing a guideline of appropriate responses, feeding in-context examples, and leveraging batch generation to save on costs. As a first baseline, we query GPT-4 with a single prompt (denoted as $p_{\text{sim}}^{\text{GPT-4}}$) and we find $p_{\text{sim}}^{\text{GPT-4}}$ has a high agreement rate with human annotators (65%; see results in Section 4.3). However, we find that this simple baseline of $p_{\text{sim}}^{\text{GPT-4}}$ fails to capture the variability in human annotation and can lead to qualitatively different results for method development, especially for reward over-optimization (Section 4.3).
@@
To more completely emulate human annotators, we modify the basic simulated annotator design to capture annotator variability in two ways. First, we emulate inter-annotator variability in the simulated pairwise preference $p_{\text{sim}}$ by mimicking a pool of annotators. We design different annotators by querying different API LLMs and varying the prompts with different formats, batch sizes, and in-context examples. In the end, we created 13 simulated annotators which we describe fully in Appendix C. Second, we emulate intra-annotator variability by directly injecting random noise and flipping the simulated preference $25\%$ of the time.
@@ 3.3 Designing an automatic evaluation
To quantify the quality of an LLM $p_{\theta}$, we measure the win-rate of that LLM against a reference model, i.e, the expected number of times that the output from $p_{\theta}$ is preferred to the output of a reference model $p_{\text{ref}}$ on the same instruction $x$. The benefits of using simulated win-rates are that it provides a metric that is easy to understand, is comparable across methods conditioned on a single reference model, and can reuse the routine we built for pairwise feedback. We use the 13 simulated annotators described in Section 3.2 without injecting the additional noise (as adding uniform noise does not change model rankings) and denote this preference simulator as $p_{\text{sim}}^{\text{eval}}$. For the reference model, we use Davinci003 as it is a well-studied system that performs similarly to the models we fine-tune.
Evaluation data.
Instruction following requires diverse coverage over realistic interactions. To build an appropriate evaluation protocol, we combine several open-source evaluation datasets and use real-world interactions with a demo instruction-following LM (Alpaca Demo (Taori23, )) as guidance for constructing our data combination. Due to privacy concerns, we do not directly release the demo data and opt to use it to guide how we combine existing open evaluation datasets.
Our final evaluation dataset consists of 805 instructions, which includes 252 instructions from the self-instruct evaluation set (Wang22*c, ), 188 from the Open Assistant (OASST) evaluation, 129 from the helpful evaluation released by Anthropic (Bai22*a, ), 80 from Vicuna evaluation (zheng2023judging, ; Chiang23*a, ), and 156 from Koala evaluation (Geng23, ). In Table 1 and Figure 2, we show example instructions from our evaluation dataset and their root verb distribution, which shows the diverse coverage of our evaluation data. We find that aggregating across datasets is important for automatic evaluation to match real-world interactions, as discussed in Section 4.4.
@@ 3.4 Reference methods in AlpacaFarm
Finally, AlpacaFarm defines a collection of validated LPF methods for instruction following. We leave a more detailed methods description to Appendix A and provide a brief overview here. In all LPF methods that follow, we begin by first performing an initial fine-tuning step on supervised data of instructions and outputs.
We begin by describing three baselines that directly operate on pairwise feedback.
@@
  • Best-of-$n$ sampling. Best-of-$n$ (or re-ranking) (Stiennon20, ; Askell21, ; Glaese22, ; Bakker22, ) is a simple but effective inference-time method that draws $n$ i.i.d. responses from the SFT model and returns the response with the highest surrogate reward.
@@
  • Proximal Policy Optimization (PPO). PPO (Kakade02, ; Schulman17, ) is a popular reinforcement learning algorithm that maximizes surrogate reward, subject to a KL penalty keeping parameters near SFT initialization.
  • Quark. We use the top-quantile variant of Quark Lu22*b which bins sequences by reward and trains on the best bin, along with adding KL and entropy regularization.

### 08 Jan 2024 v4

@@ Abstract
Large language models (LLMs) such as ChatGPT have seen widespread adoption due to their strong instruction following abilities. Developing these LLMs involves a complex yet poorly understood workflow requiring training with human feedback. Replicating and understanding this instruction-following requires tackling three major challenges: the high cost of data collection, the lack of trustworthy evaluation, and the absence of reference method implementations. We address these challenges with AlpacaFarm, a simulator that enables research and development for learning from feedback at a low cost. First, we design LLM prompts to simulate human feedback that are 50x cheaper than crowdworkers and display high agreement with humans. Second, we propose an automatic evaluation and validate it against human instructions obtained on real-world interactions. Third, we contribute reference implementations for several methods (PPO, best-of-$n$, expert iteration, and more) that learn from pairwise feedback. Finally, as an end-to-end validation of AlpacaFarm, we train and evaluate eleven models on 10k pairs of real human feedback and show that the rankings of models trained in AlpacaFarm match the rankings of models trained on human data. As a demonstration of the research possible in AlpacaFarm, we find that methods that use a reward model can substantially improve over supervised fine-tuning and that our reference PPO implementation leads to a +10% improvement in win-rate against Davinci003. We release all components of AlpacaFarm at https://github.com/tatsu-lab/alpaca_farm.
Figure 1: AlpacaFarm is a simulation sandbox that enables fast and cheap experimentation on LLMs that learn from human feedback. It simulates human feedback with (oracle) API LLMs, provides a validated evaluation protocol, and offers a suite of reference method implementations. Researchers can rapidly iterate on model development and transfer their methods to training on human data to maximize performance.
@@ 1 Introduction
Large language models (LLMs) (Bommasani21, ; Brown20, ; OpenAI23, ) have demonstrated unprecedented capabilities in following diverse and open-ended instructions (Ouyang22, ; Askell21, ; Longpre23, ). These achievements have often been attributed to the fine-tuning of pretrained LLMs using human feedback, but this process remains poorly understood due to the lack of published information on the training methods from LLM vendors. For example, it was recently revealed that only the Davinci003 model in the instruct series of OpenAI models used reinforcement learning (RL) with the PPO algorithm OpenAI*a , leading some to question the importance of RL in the training process given that Davinci002 already worked very well. Understanding and improving these methods requires open and transparent replications of the training process, but this remains challenging due to the cost and complexity associated with methods that learn from human feedback.
@@
Our goal is to facilitate research and development on instruction following models and methods that learn from human feedback. We identify three main challenges: the high cost of data annotation, the lack of automated evaluation for model development, and the absence of validated implementations of existing methods. To address these three challenges, we introduce AlpacaFarm (Figure 1), a simulation sandbox that enables experimentation at low cost. Using AlpacaFarm, researchers can rapidly iterate on method development in simulation and transfer these insights to build high-performance systems with actual human feedback.
@@
For the first challenge of data annotation costs, AlpacaFarm simulates human annotators with (oracle) API LLMs that are faster and cheaper. To collect simulated feedback data, we design prompts for oracle API LLMs (e.g. GPT-4) that enable us to simulate human pairwise comparisons at a cost that is 45x cheaper than crowdworkers, and tune these prompts to faithfully capture many aspects of human annotators, such as their quality judgments, inter-annotator variability, and stylistic preferences.
@@
For the second challenge of automated evaluation, we designed an automatic evaluation protocol aiming to quantify system performance on simple yet realistic real-world human instructions. Improving evaluations of open-ended instruction following has been challenging due to the cost and non-replicability of human evaluation, the lack of real human interaction data, and the diversity of natural human instructions. To address this, we used the aforementioned simulated annotators to compute the win-rate between outputs from the evaluated model and those from a baseline LLM on a set of simple yet realistic instructions. To select the instructions, we used a set of real-world user interactions (Alpaca Demo (Taori23, )) as reference instructions. As those cannot be released directly due to confidentiality constraints, we combined existing public evaluation datasets to create a set of instructions that mimic the Alpaca Demo evaluation set. Quantitative evaluations of system rankings on our evaluation data show a high correlation with system rankings on the Alpaca Demo instructions.
@@
For the third challenge of missing reference implementations, we implement and test several popular learning algorithms including PPO (Schulman17, ), expert iteration (Anthony17, ), and Quark Lu22*b , and release reference implementations. We show that, among the methods we studied, PPO with a surrogate reward model is the most effective training-time method, improving the win-rate against Davinci003 of an instruction fine-tuned LLaMA 7B model from 44% to 55%. Other baselines that have been validated on simpler tasks fall short in comparison, highlighting the importance of testing these algorithms in a real instruction-following setting.
@@
As an end-to-end evaluation of the AlpacaFarm, we compare eleven methods trained and evaluated in AlpacaFarm with the same methods trained and evaluated on actual human feedback. We show that the method rankings obtained from developing on AlpacaFarm closely agree with the method rankings obtained from training on actual human data (Spearman correlation of 0.98) and that the best method in AlpacaFarm leads to substantial gains with human feedback. Finally, we find that AlpacaFarm can replicate qualitative behaviors of human feedback such as over-optimization of the reward model, suggesting that AlpacaFarm serves as an effective way for researchers to rapidly study and develop methods for training instruction-following LLMs using human feedback.
@@ 2 Background & problem statement
To begin, Section 2.1 introduces the instruction following task and the pairwise-comparison-based human feedback setting that we study. With this background, we formally define in section 2.2 the goals of developing a low-cost simulator for studying instruction following and learning from human feedback.
@@ 2.1 Learning to follow instructions
While there are a rich set of methods that learn from human feedback to directly optimize $R$ (see Section 6), in this work we focus on the setting of learning from pairwise feedback (LPF) due to its central role in recent instruction-following LLMs Ouyang22 . The starting point of this process is a model that is fine-tuned on instruction-following demonstrations $(x,y)$, which we denote as $p^{\text{SFT}}_{\theta}(y\mid x)$. The LPF process then involves taking pairs of samples $y_{0},y_{1}$ from $p^{\text{SFT}}_{\theta}$ for each instruction $x$, querying humans for which sample within each pair is better, and learning from this pairwise feedback. Since all methods start from the SFT base, we use $p_{\theta}$ for notational simplicity.
@@
Once instruction-following models are trained, researchers need to evaluate these models. One common approach for evaluating models is pairwise model evaluation (Chiang23, ; Geng23, ; Bai22, ), which performs pairwise comparisons on outputs generated by the model $p_{\theta}$ and a reference model $p_{\text{ref}}$ similarly to the pairwise feedback process. Concretely, we collect pairwise preference for two models ($\{(x^{(j)},y_{\theta}^{(j)},y_{\text{ref}}^{(j)},z^{(j)})\}_{j}$), which is aggregated by computing the average win-rate – the percentage of times $p_{\theta}$ is preferred to $p_{\text{ref}}$. Researchers can then compare LPF models by their win-rates against the same reference model.
@@ 2.2 Problem statement
For pairwise feedback, we substitute human preference judgements $z_{\text{human}}\sim p_{\text{human}}(z\mid x,y_{0},y_{1})$ with a simulated preference $z_{\text{sim}}\sim p_{\text{sim}}(z\mid x,y_{0},y_{1})$ using oracle API LLMs. Our goal is to construct a $p_{\text{sim}}$ that is both low-cost and faithfully captures different aspects of human preference feedback, such as quality judgments, inter-annotator agreement rates, and stylistic preferences.
For evaluations, we evaluate system outputs using the pairwise preference simulator and identify evaluation datasets that reflect natural human-LLM interactions. The goal for our evaluations is to ensure that system rankings on the new evaluation dataset closely match human rankings on a realistic set of instructions. We use the Alpaca Demo Ouyang22 as reference for real-world instructions.
@@
AlpacaFarm combines these three components into a simulation framework for learning from pairwise feedback. We evaluate the complete system by an end-to-end workflow of developing methods in simulation and transferring the insights to the real world. Concretely, we will run each method $M$ on the simulated preferences (called $M_{\text{sim}}$) and evaluate with simulated rankings $p_{\text{sim}}$. In parallel, we will run $M$ on human preferences (called $M_{\text{human}}$) and evaluate with human rankings $p_{\text{human}}$. We consider AlpacaFarm to be successful if the simulated method rankings correlate well with the human method rankings. Note that the goal of AlpacaFarm is not to train a good model with simulated data, instead it is to analyze methods, i.e., algorithms and hyperparameters, with simulated data and transfer the insights to the real world.
The rest of this work will present the details of the pairwise feedback and evaluation design (Section 3), evaluate these designs (Section 4), and analyze the different reference methods we implemented in the AlpacaFarm (Section 5).
@@ 3.2 Designing simulated pairwise preference $p_{\text{sim}}$
Equipped with the Alpaca instruction data, we now describe the design of our simulated annotator for pairwise preferences. Our core proposal is to design the simulator $p_{\text{sim}}(z\mid x,y_{0},y_{1})$ by prompting OpenAI API LLMs. While using LLMs as a proxy for annotators has become increasingly popular (Chiang23, ; Liu23*a, ), using LLMs as part of a simulation environment poses major additional challenges. Our simulated preferences $p_{\text{sim}}$ must not only have a high agreement with human preferences $p_{\text{human}}$, it must also capture other qualitative aspects of human feedback such as inter- and intra-annotator inconsistencies. Intuitively, the noise and variability in pairwise feedback are key parts of the challenge in the LPF problem and we find that ignoring these factors leads to a simulator that diverges substantially from real-world behavior (Section 4.3). Indeed, when training on sampled preference the variability of the preference may affect the LPF training dynamics.
@@
To start with, we design prompts by providing a guideline of appropriate responses, feeding in-context examples, and leveraging batch generation to save on costs. As a first baseline, we query GPT-4 with a single prompt (denoted as $p_{\text{sim}}^{\text{GPT-4}}$) and we find $p_{\text{sim}}^{\text{GPT-4}}$ has an agreement rate with human annotators that is similar to the agreement rate between humans (65% vs 66%; see results in Section 4.3). However, we find that this simple baseline of $p_{\text{sim}}^{\text{GPT-4}}$ fails to capture the variability in human annotation and can lead to qualitatively different results for method development, especially for reward over-optimization (Section 4.3).
@@
To better emulate human annotators, we modify the basic simulated annotator design to capture annotator variability in two ways. First, we emulate inter-annotator variability in the simulated pairwise preference $p_{\text{sim}}$ by mimicking a pool of annotators. We design different annotators by querying different API LLMs and varying the prompts with different formats, batch sizes, and in-context examples. In the end, we created 13 simulated annotators which we describe fully in Appendix C. Second, we emulate intra-annotator variability by flipping the simulated preference $25\%$ of the time, i.e., we inject random noise.
@@ 3.3 Designing an automatic evaluation
To quantify the quality of an LLM $p_{\theta}$, we measure the win-rate of that LLM against a reference model, i.e, the expected number of times that the output from $p_{\theta}$ is preferred to the output of a reference model $p_{\text{ref}}$ on some instruction $x$. The benefits of using simulated win-rates are that it provides a metric that is easy to understand, is comparable across methods conditioned on a single reference model, and can reuse the routine we built for pairwise feedback. We use the 13 simulated annotators described in Section 3.2 without injecting the additional noise (as adding uniform noise does not change model rankings) and denote this preference simulator as $p_{\text{sim}}^{\text{eval}}$. For the reference model, we use Davinci003 as it is a well-studied system that performs similarly to the models we fine-tune.
Evaluation instructions.
Instruction following requires diverse coverage over realistic interactions. To build an appropriate evaluation protocol, we combine several open-source evaluation datasets and use real-world interactions with a demo instruction-following LM (Alpaca Demo (Taori23, )) as guidance for constructing our data combination. Due to privacy considerations, we do not directly release the demo data and opt to use it to guide how we combine existing open evaluation datasets.
We construct an evaluation of 805 instructions, which includes 252 instructions from the self-instruct test set (Wang22*c, ), 188 from the Open Assistant (OASST) test set, 129 from Anthropic’s helpful test set (Bai22*a, ), 80 from Vicuna test set (zheng2023judging, ; Chiang23*a, ), and 156 from Koala test set (Geng23, ). In Table 1 and Figure 2, we show example instructions from our evaluation dataset and their root verb distribution, which shows the diverse coverage of our evaluation data. We find that combining across datasets is important for automatic evaluation to match real-world interactions, as discussed in Section 4.4.
@@ 3.4 Reference methods in AlpacaFarm
Finally, AlpacaFarm defines a collection of validated LPF methods for instruction following. We leave a more detailed methods description to Appendix A and provide a brief overview here. In all LPF methods that follow, we begin by first performing an initial fine-tuning step on supervised data consisting of instructions and outputs. We begin by describing three baselines that directly operate on pairwise feedback.
@@
  • Best-of-$n$ sampling. Best-of-$n$ (or re-ranking) (Stiennon20, ; Askell21, ; Glaese22, ; Bakker22, ) is a simple but effective inference-time method that draws $n$ i.i.d. responses from the SFT model and returns the response with the highest surrogate reward. Unless stated otherwise, we use $n=1024$ in our experiments.
@@
  • Proximal Policy Optimization (PPO). PPO (Kakade02, ; Schulman17, ) is a popular reinforcement learning algorithm that maximizes surrogate reward, subject to a KL penalty keeping parameters near the SFT initialization.
  • Quark. We use the top-quantile variant of Quark Lu22*b , which bins examples by reward and trains on the best bin, while minimizing the KL divergence between the trained model and the SFT initialization, and maximizing the model’s entropy.